# How Does Stake Distribution Influence Consensus? Analyzing Blockchain Decentralization

### v1

@@ I Introduction
Bitcoin shaped the field of blockchains by introducing a peer-to-peer system that operates without trusted intermediaries [70]. Its vision encapsulates the essence of decentralization, characterized by the elimination of single points of failure and the facilitation of collective decision-making. Our work focuses on exploring decentralization in blockchains, particularly in their consensus mechanisms. These mechanisms are critical as they establish agreement on the content and order of transactions among validators.
@@ II-A Classification of Consensus Mechanisms Based on Finality
Probabilistic Finality is achieved when the finality of an appended block $b$ at time $t_{0}$ is expressed as $f(b,t_{0})=1-\gamma$, where $\gamma$ is a non-negative value less than 1 (i.e., $0\leq\gamma<1$) that represents the deviation from total finality. Specifically, for any two points in time $t_{1}$ and $t_{2}$ such that $t_{2}>t_{1}$, it follows that $f(b,t_{1})<f(b,t_{2})$. As time progresses, $f(b,t)$ gradually converges to 1 as $\gamma$ approaches zero.
$\lim_{\gamma\to 0}f(b,t)=1\vskip-6.0pt$ (1)
@@ II-B Transition to Weighted Consensus
$\mathbb{Q}\geq\left(\frac{2}{3}\right)m\text{, where }m=|N|\vskip-2.0pt$ (2)
@@
$\mathbb{Q}^{\prime}\geq\left(\frac{2}{3}\right)\sum_{k}w_{k}\quad\forall k\in N\vskip-2.0pt$ (4)
@@ III-3 Nakamoto Coefficient - Liveness ($\mathbb{N}_{L},\rho_{\mathbb{N}_{L}}$)
The normalized Nakamoto coefficient $\rho_{\mathbb{N}_{L}}$ is then calculated as:
$\rho_{\mathbb{N}_{L}}=\frac{\mathbb{N}_{L}}{m}$ (7)
@@
A higher $\mathbb{N}_{L}$ suggests better decentralization. The normalized value $\rho_{\mathbb{N}_{L}}$, ranging between 0 and 1. A $\rho_{\mathbb{N}_{L}}$ closer to 1 indicates high decentralization and resilience against censorship [32].
@@ III-4 Nakamoto Coefficient - Safety ($\mathbb{N}_{S},\rho_{\mathbb{N}_{S}}$)
The normalized form of $\mathbb{N}_{S}$, denoted by $\rho_{\mathbb{N}_{S}}$, is:
@@
Higher values of $\mathbb{N}_{S}$ and $\rho_{\mathbb{N}_{S}}$ indicate a more decentralized system. Specifically, $\rho_{\mathbb{N}_{S}}$ close to 1, especially in systems with a large validator set ($m$), signifies robust decentralization.
@@ IV Empirical Data Analysis
TABLE III: Decentralization metrics for blockchains, as of 14 December 2023
@@
Aptos L1 blockchain [7] HotStuff/DiemBFT [26] 144 0.56 12.50 (18) 12.50 (38) 8.488454e+11 7.63
@@ IV-A Scope and Methodology of Data Collection
In our study, we focus on permissionless blockchains that use classical consensus mechanisms, particularly those with weighted consensus as outlined in Section II. We limit our analysis to blockchains with deterministic validator set selection methods, such as PoS and DPoS. Notably, all the blockchains examined in this study employ DPoS for Sybil attack resistance.
For our empirical analysis, we selected ten blockchains, as shown in Table III, namely: Aptos, Axelar, BNB (Binance), Celestia, Celo, Cosmos, Injective, Osmosis, Polygon, and Sui. These protocols are based on BullShark [82], HotStuff [93], IstanbulBFT [64], and Tendermint [14] consensus mechanisms. These blockchains were chosen for their diverse applications, including smart contracts Layer-1 (L1) blockchains, interoperability protocols [11], Layer-2 (L2) scaling solutions [59], data availability protocols [5], and decentralized exchanges [50], ensuring the wide applicability of our findings. Our sample comprises blockchains with at least 300 million USD market capitalization, cumulatively amounting to a market capitalization of 60 billion USD as of December 14, 2023 [1].
@@ IV-B Data Analysis
Challenge 3: How to discourage the creation of multiple Sybil identities?
@@ V-A Primitives and Assumptions
[…]
Capping the validator set cardinality is justified for two reasons. Firstly, in line with classical consensus mechanisms, an increase in the number of validators tends to decrease the system scalability, measured in throughput and latency [93]. Secondly, by imposing a maximum limit on the number of validators, and a minimum capital requirement of $s_{M}$ staked tokens for running a validator, the system discourages single entities from dominating the validator set, a point further explored in the following section.
@@ V-B SRSW Function
$\mathbb{Q}^{*}\geq\left(\frac{2}{3}\right)\sum_{i}w_{i}^{*}=\left(\frac{2}{3}\right)\sum_{i}\sqrt{s_{i}}\quad\forall i\in N\vskip-4.0pt$ (13)
@@
$r^{*}_{n_{i}}=\alpha w_{i}^{*}=\alpha\sqrt{s_{i}},\text{ if }s_{i}>s_{M};\text{ otherwise, }0.$ (14)
@@
$r_{n_{i}}^{*}>r_{n_{j}}^{*}+r_{n_{k}}^{*}-C,\text{ where }s_{i}\geq s_{j}+s_{k}\vskip-4.0pt$ (15)
@@ VI-A Decentralization Metrics Analysis
$G$ % decrease $\mathbb{N}_{L}\%$ increase $\mathbb{N}_{S}\%$ increase
@@ Acknowledgment
This work has in part been supported by NSERC.

### v3

@@ I Introduction
Bitcoin shaped the field of blockchains by introducing a peer-to-peer system that operates without trusted intermediaries [71]. Its vision encapsulates decentralization, characterized by the elimination of single points of failure and the facilitation of collective decision-making. Our work focuses on exploring decentralization in blockchains, particularly in their consensus mechanisms. These mechanisms are critical as they establish agreement on the content and order of transactions among validators.
@@ II-A Classification of Consensus Mechanisms Based on Finality
Probabilistic Finality is achieved when the finality of an appended block $b$ at time $t_{0}$ is expressed as $f(b,t_{0})=1-\gamma$, where $\gamma$ is a non-negative value less than 1 (i.e., $10\leq\gamma<$ 1) that represents the deviation from total finality. Specifically, for any two points in time $t_{1}$ and $t_{2}$ such that $t_{2}>t_{1}$, it follows that $f(b,t_{1})<f(b,t_{2})$. As time progresses, $f(b,t)$ gradually converges to 1 as $\gamma$ approaches zero.
$\lim_{\gamma\to 0}f(b,t)=1\vspace{-6pt}$ (1)
@@ II-B Transition to Weighted Consensus
$\mathbb{Q}\geq\left(\frac{2}{3}\right)m\text{, where }m=|N|\vspace{-2pt}$ (2)
@@
$\mathbb{Q}^{\prime}\geq\left(\frac{2}{3}\right)\sum_{k}w_{k}\quad\forall k\in N\vspace{-2pt}$ (4)
@@ III-3 Nakamoto Coefficient - Liveness ($\mathbb{N}_{L},\rho_{\mathbb{N}_{L}}$)
The scaled Nakamoto coefficient $\rho_{\mathbb{N}_{L}}$ is then calculated as:
$\rho_{\mathbb{N}_{L}}=\frac{\mathbb{N}_{L}}{m}*100$ (7)
@@
A higher $\mathbb{N}_{L}$ suggests better decentralization. The normalized value $\rho_{\mathbb{N}_{L}}$, ranging between 0 and 100. A $\rho_{\mathbb{N}_{L}}$ closer to 100 indicates high decentralization and resilience against censorship [32].
@@ III-4 Nakamoto Coefficient - Safety ($\mathbb{N}_{S},\rho_{\mathbb{N}_{S}}$)
The scaled form of $\mathbb{N}_{S}$, denoted by $\rho_{\mathbb{N}_{S}}$, is:
@@
Higher values of $\mathbb{N}_{S}$ and $\rho_{\mathbb{N}_{S}}$ indicate a more decentralized system. Specifically, $\rho_{\mathbb{N}_{S}}$ close to 100, especially in systems with a large validator set ($m$), signifies robust decentralization.
@@ IV Empirical Data Analysis
TABLE III: Decentralization metrics for blockchains, as of 14 December 2023. Ideal values of G and $\varepsilon$ are 0.
@@
Aptos L1 blockchain [7] Jolteon/DiemBFT [26, 37] 144 0.56 12.50 (18) 12.50 (38) 8.488454e+11 7.63
@@ IV-A Scope and Methodology of Data Collection
In our study, we focus on permissionless blockchains that use classical consensus mechanisms, particularly those with weighted consensus as outlined in Section II. We limit our analysis to blockchains with deterministic validator set selection methods, such as PoS and DPoS. Notably, all the blockchains examined in this study employ DPoS for Sybil resistance.
For our empirical analysis, we selected ten blockchains, as shown in Table III, namely: Aptos, Axelar, BNB (Binance), Celestia, Celo, Cosmos, Injective, Osmosis, Polygon, and Sui. These protocols are based on BullShark [83], HotStuff [94], IstanbulBFT [65], and Tendermint/CometBFT [14] consensus mechanisms. These blockchains were chosen for their diverse applications, including smart contracts Layer-1 (L1) blockchains, interoperability protocols [11], Layer-2 (L2) scaling solutions [60], data availability protocols [5], and decentralized exchanges [51], ensuring the wide applicability of our findings. Our sample comprises blockchains with at least 300 million USD market capitalization, cumulatively amounting to a market capitalization of 60 billion USD as of December 14, 2023 [1].
@@ IV-B Data Analysis
Challenge 3: How to discourage creation of multiple Sybil identities?
@@ V-A Primitives and Assumptions
$r_{n_{k}}=\begin{cases}\alpha w_{k}\quad&\text{if}\,s_{k}>s_{M},\\ 0\quad&\text{if}\,s_{k}\leq s_{M}.\\ \end{cases}$ (11)
Capping the validator set cardinality is justified for two reasons. Firstly, in line with classical consensus mechanisms, an increase in the number of validators tends to decrease the system scalability, measured in throughput and latency [94]. Secondly, by imposing a maximum limit on the number of validators, and a minimum capital requirement of $s_{M}$ staked tokens for running a validator, the system discourages single entities from dominating the validator set (i.e., Sybil attacks), further explored in the following section.
@@ V-B SRSW Function
$\mathbb{Q}^{*}\geq\left(\frac{2}{3}\right)\sum_{i}w_{i}^{*}=\left(\frac{2}{3}\right)\sum_{i}\sqrt{s_{i}}\quad\forall i\in N\vspace{-4pt}$ (13)
@@
[…]
@@
$r_{n_{i}}^{*}>r_{n_{j}}^{*}+r_{n_{k}}^{*}-C,\text{ where }s_{i}\geq s_{j}+s_{k}\vspace{-4pt}$ (15)
@@ VI-A Decentralization Metrics Analysis
$G$ decrease $\mathbb{N}_{L}\%$ increase $\mathbb{N}_{S}\%$ increase
@@ Acknowledgment
This work has in part been supported by Sui Foundation and NSERC.